OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x^2 + 8x + 18

Parth (parthkohli):

Do you know how to complete the square?

OpenStudy (anonymous):

(b/2)^2

Parth (parthkohli):

Yes, so can you apply that here? I'm sure you can...

OpenStudy (anonymous):

here are the choices: y = (x - 4)^2 - 2 y = (x - 4)^2 + 2 y = (x + 4)^2 - 2 y = (x + 4)^2 + 2

Parth (parthkohli):

Okay, but can you add \(\left(b\over 2\right)^2\) to both sides? Do you know what \(b\) is?

OpenStudy (anonymous):

8

Parth (parthkohli):

So what would \(\left(b\over 2\right)^2\) be?

OpenStudy (anonymous):

16?

Parth (parthkohli):

Yes, so add \(16\) to both sides. :-)

OpenStudy (anonymous):

y = (x + 4)^2 - 2

Parth (parthkohli):

Please show your work.